[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Oct 13, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Oct 13 17:42:34 UTC 2016


HSU Shift Log for Thu Oct 13 2016 13:42:29 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues: Issue retrieving SHIPS guidance at 12z due to the  
supercomputer switch. Had to send and retrieve guidance a second time which  
resolved the problem.

N-AWIPS/AWIPS2 issues:

Other Techincal issues: Marc Laurent tried fixing the Java script issue  
with the Bermuda webpage. It seemed to be working at first, but after the  
PCs were rebooted, the radar/data would not display again. Radar data and  
some of the Bermuda observations can be accessed from a separate part of  
the website in which Java is not an issue:  
http://www.weather.bm/tools/graphics.asp

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161013/12c39833/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list